OpenStudy (anonymous):

Complete the conversion statement. 0.7 mm = ____ cm A. 70.0 B. 0.07 C. 7.0 D. 700

OpenStudy (anonymous):

B

OpenStudy (anonymous):

@iGreen may you please double check his/her answer please

OpenStudy (igreen):

He's correct, just move the decimal place over to the right.

OpenStudy (anonymous):

ok mind helping with some more

OpenStudy (anonymous):

please

OpenStudy (anonymous):

heres the link to the question

OpenStudy (anonymous):

: The perimeter of this rectangle can be calculated using the equation P = 2 × 18.5 + 2 × 7.2. Which other equation could be used to calculate the perimeter? A. P = 2 × 18.5 + 7.2 B. P = 2 (18.5 + 7.2) C. P = 2 × 18.5 × 7.2 D. P = (2 • 18.5) + 7.2

OpenStudy (anonymous):

@iGreen help help help me please